if someone had a batting average of .311 during a regular season, how many hits would you expect him to get in the playoffs if he has a total or 68 at-bats?
Answer by stanbon(75887)

if someone had a batting average of .311 during a regular season, how many hits would you expect him to get in the playoffs if he has a total or 68 at-bats?
-------
x/68 = 313/1000
---
x = 68(0.313)
---
x = 21.28 hits
======================
Cheers,
Stan H.